# Transformation equation for frames undergoing non-uniform acceleration such as SHM and rotational motion

### v2

@@ Abstract
The Lorentz transformation equations provide us a set of relations between the spacetime coordinates as observed from two different inertial frames. In case one of the frames is moving with a uniform rectilinear acceleration, we have Rindler’s transformation equations under such a situation. In the present work, we extend the Rindler’s equations to a situation where we generally have non-uniform acceleration. After that, we consider the non-inertial frame to undergo simple harmonic motion (SHM), and as a second case, we consider the non-inertial frame to move uniformly along a circle. This set of transformation equations will have applications in various branches of Physics and in general in Astrophysics.
Keywords: Relativity; Frame of references; Rindler transformation; Non-uniform motions; Circular motions.
@@ 1 Introduction
In special theory of relativity, the coordinate transformation relations for a frame of reference moving with uniform velocity with respect to a rest frame are known as the Lorentz transformation [1]. These relations play a vital role during the investigation of a test particle in a frame moving with a uniform velocity. In the case of an accelerated frame, the general transformation relations are presented by Nelson [2, 3]. The coordinate transformations for uniformly accelerated frame with constant acceleration is known as the Rindler transformation, and the consequences of such a coordinate frame have been studied by different researchers [4, 5, 6, 7, 8, 9]. In the Rindler transformation, the initial velocity is considered as zero. However, the coordinate transformation for an accelerated frame with non-zero initial velocity has been discussed by Yi [6]. Additionally, Brahma and Sen [10] analyzed the Rindler transformation having acceleration along an arbitrary direction on the XY plane. In the present work, we investigate the coordinate transformation relations for a non-uniform acceleration such as a frame executing simple harmonic motion (SHM) and subsequently the transformation relations for a frame having uniform circular motion (UCM). However, in special relativity, when the motion of a test particle is subject to a restoring force along the $X$-direction, the mass of the particle can not be considered as constant. But it changes according to the law of special theory of relativity, to which MacColl [11] calculated a general expression of the time period of the relativistic oscillation, which is a function of the total energy of the system. Such a relativistic harmonic oscillation can be used to admit a consistent procedure for the quantization of the system [12]. However, the formal investigation on the harmonic oscillator in relativity using the Lagrange formalism [13] and the Hamiltonian formalism [14] indicates that even though the term “harmonic oscillator” has been generally used, but the results obtained during the calculations are no longer harmonic in nature. Instead, it is “anharmonic” in nature, which is due to the relativistic effect on mass and energy. It indicates that if the maximum velocity of the SHM is comparable to the speed of light, then the harmonic oscillation becomes an anharmonic in nature. The nature of the relativistic harmonic oscillation and the associated anharmonicity due to the relativistic effect have been experimentally studied using ultracold lithium atoms in the third band of an optical lattice and reported by Fujiwara et al. [15]. Here, it has been stated that the measured worldline shapes, relativistic anharmonicity, monopole oscillations, and relativistic dephasing of oscillator ensembles are in good agreement with the theoretical predictions of the relativistic harmonic oscillation. It is to be mentioned that, although the simple harmonic oscillation has been studied extensively in the regime of relativity but, only a few works on relativistic coordinate transformation between an inertial frame and a frame executing SHM have been reported in literature, to our knowledge. For example, Perepelkin et al. [16] discussed the transformation relation for a frame having velocity which is a sinusoidal function of time by using the generalized Galilean transformation. Therefore, it would be very much important and appropriate, to study the transformation relations for the frame undergoing SHM through a more general procedure. This motivated us to calculate such coordinate transformation relations in the present work.
@@
So in the present context, in order to carry out the calculations to obtain the coordinate transformations between an inertial frame and a frame executing SHM, we have to first derive Rindler like transformation equations for coordinates frames having non-uniform acceleration. Since the circular motion of an object can be described by two SHMs with same amplitudes in perpendicular directions having phase difference of $\pi/2$ radians so, the transformation relations for the frame executing SHM can be used to understand the dynamics of a moving frame in a circular path. As a result, we obtain the coordinate transformation relations for a frame of reference moving in a circular path. The results in the present work will have immense applications in the study of the test particles in non-uniformly accelerated frames such as frames executing SHM and the frames moving in a circular path.
@@
We organized our paper as: In Section 2 we obtain the transformation relations for a frame of reference executing simple harmonic motion (SHM) along the $X$-direction. Additionally, we also calculated the similar transformation relations for the frame of reference executing SHM along the $Y$-direction with initial phase of $\pi/2$ radians. And then in Section 3, as mentioned above since the circular motion can be described by using the two SHMs in perpendicular direction and phase difference of $\pi/2$ radians, towards the end we obtained the transformation relations for a frame that is moving in a circular path. In Section 4 we compared our results with the existing coordinate transformation relations for the frames executing SHM and UCM available in literature. At last, in Section 5 we discussed the results obtained in the above calculations and concluded our findings.
@@ 2 Transformation relations for a frame of reference executing SHM
The original Rindler transformation equation is meant for constant acceleration. However, in the present work, we shall derive the transformation relations in which proper acceleration or coordinate acceleration is a function of time. Most specifically for a frame executing simple harmonic motion (SHM) the coordinate acceleration is a sinusoidal function of time. Below, we derive the transformation relations for the case when the frame is oscillating along the $X$-direction having displacement proportional to some sinusoidal function of angular frequency $\omega$.
@@
where $u_{x}=\frac{dx}{dt}$, $a_{x}=\frac{du_{x}}{dt}$ and $a^{\prime}_{x}=\frac{du^{\prime}_{x}}{dt^{\prime}}$. Therefore, for a particle that has acceleration $a^{\prime}_{x}$ with respect to an inertial frame $S^{\prime}$ which instantaneously accompanies the particle $(u_{x}=v)$, we can write:
@@
In case of Castillo & Sanchez [8] they assumed that the acceleration $a^{\prime}_{x}$ is constant and obtain the Rindler’s transformation equation [17]. However, in our present case we shall consider $a^{\prime}_{x}$ to be a function of $(x^{\prime},t^{\prime})$ and proceed as follows.
@@
The quantity $u_{x}$ in the right hand side is actually representing $v$. So we can identify $\left(1-\frac{u_{x}^{2}}{c^{2}}\right)^{1/2}dt=d\tau$, where $\tau$ is the proper time. Also, as mentioned by Castillo & Sanchez [8] the time $t^{\prime}$ as measured in $S^{\prime}$ frame is actually the proper time $\tau$. So we can finally write:
@@
Integrating on both side,
@@
Equations (6) and (7) with appropriate constants of integration will help us to find appropriate co-ordinate transformation equations where the acceleration is not constant, but varying with time.
However, in our present case, we take $a^{\prime}_{x}$ to be a function of timetime and we shall consider two standard cases: (i) the frame $S^{\prime}$ undergoing Simple Harmonic Motion (SHM) and (ii) the frame $S^{\prime}$ under going Circular Motion.
@@ 2.1 The frame $S^{\prime}$ undergoing SHM in $X$-direction:
In the first case, we consider that acceleration $a^{\prime}_{x}$ is a function of proper time $\tau$ or $(t^{\prime})$ and represents a simple harmonic motion (SHM). So we may take a displacement $x^{\prime}=r_{0}\sin(\omega\tau)$ such that the velocity of the oscillation $u^{\prime}_{x}=\frac{dx^{\prime}}{d\tau}=r_{0}\omega\cos(\omega\tau)$ and the acceleration $a^{\prime}_{x}=\frac{d^{2}x^{\prime}}{d\tau^{2}}=-r_{0}\omega^{2}\sin(\omega\tau)$, where $r_{0}$ is the amplitude of the SHM and $\omega$ is the angular frequency. Therefore, substituting $a_{x}^{\prime}=-r_{0}\omega^{2}\sin(\omega\tau)$ in equation (4) we obtain:
@@
Now integrating both side of the equation (9) we obtain as:
@@
The right hand side (RHS) of the above equation (13) is analytically not integrable. However, as in this equation $r_{0}\omega$ represents the maximum value of the linear velocity of the frame executing simple harmonic motion (SHM), and for practical purposes this velocity will be much smaller than the speed of light $(c)$. Thus assuming $\frac{r_{0}\,\omega}{c}<<1$, we can write: $\frac{r_{0}\omega}{c}\cos(\omega\tau)\sim 0$ and hence the right hand side (RHS) of equation (13) can be expand in a series in the power of $r_{0}\omega/c$. Now expanding the integrand in the RHS of equation (13) and neglecting the terms higher than $(r_{0}\omega/c)^{3}$ in order, we can obtain from the above equation (13) as below:
@@
Again, since as stated above, $r_{0}\omega/c<<1$ for all practical purposes, we have $\frac{r_{0}\omega}{c}\cos(\omega\tau)\sim 0$, and hence we can take an approximation by neglecting the terms higher than $(r_{0}\omega/c)^{3}$ as before. So as before, expanding the integrand in the RHS of equation (15) in Taylor series and utilizing the above approximation we can obtain from equation (15) as:
@@
Now in the present context, as $\tau$ can be the time in the frame executing SHM, so we may replace $\tau$ by $t^{\prime}$, where $t^{\prime}$ represents the time as measured from the accelerated frame (as discussed just before equation (4)). Therefore, substituting $\tau=t^{\prime}$ we can obtain from equations (14) and (16) as follows:
@@
Then we integrate both side of equation (25), with limit 0 to $T$, for the unprimed and 0 to $T^{\prime}$, for the primed frame. i.e.
@@
Comparing (27) and (28), one can say that, we obtained $(1/4)$ as the the co-effecient of the second term in RHS, corresponding to $(3/16)$ for the same quantity as obtained by other authors [13, 15, 18].
Thus, calculations done by us are in good agreement with those approximate expressions reported by other authors in the past.

### v3

@@ Abstract
Lorentz transformation equations provide us a set of relations between the spacetime coordinates as observed from two different inertial frames. In case, one of the frames is moving with a uniform rectilinear acceleration we have Rindler’s transformation equations under such a situation. In the present work, we extend the Rindler’s equations to a situation where we have in general non-uniform acceleration. After that we consider the non-inertial frame to undergo simple harmonic motion (SHM) and as a second case we consider the non-inertial frame to move uniformly along a circle. This set of transformation equations will have applications in various branches of Physics and in general in Astrophysics.
Keywords: Relativity; Non-inertial frames; Simple harmonic motion (SHM); Uniform circular motion (UCM).
@@ 1 Introduction
In special theory of relativity, the coordinate transformation relations for a frame of reference moving with uniform velocity with respect to a rest frame are known as the Lorentz transformation [1]. These relations play a vital role during the investigation of a test particle in a frame moving with a uniform velocity. In the case of an accelerated frame, the general transformation relations are presented by Nelson [2, 3]. The coordinate transformations for uniformly accelerated frame with constant acceleration is known as Rindler transformation and the consequences of such coordinate frame have been studied by different researchers [4, 5, 6, 7, 8, 9]. In Rindler transformation the initial velocity is considered as zero. However the coordinate transformation for accelerated frame with non-zero initial velocity has been discussed by Yi [6]. Additionally, Brahma and Sen [10] analyzed the Rindler transformation having acceleration along an arbitrary direction on the XY plane. The physics with the accelerated frames can be also very efficiently handled with the help of Fermi coordinates which are local coordinates adapted to any world line ( even if it is not geodesical) [11]. Using Fermi coordinates which are basically generalization of an inertial Cartesian coordinates to an accelerated system, Nesterov [12] analyzed the plane-wave metric for a geodesic observer in a weak gravitational-wave field. In another work, Chicone and Mashhoon [13] also investigated the motion of ultra-relativistic particles and light rays using the Fermi coordinates.
Scientists also use proper reference frame in the theory of relativity, which is a particular form of accelerated reference frame where an accelerated observer can be considered as being at rest. Kajari et al. [14] discussed the propagation of light using proper reference frame. Using the hypothesis of locality, Mashhoon and Muench [15] investigated the limitations of length measurements by accelerated observers in Minkowski spacetime brought about via the hypothesis of locality, namely, the assumption that an accelerated observer at each instant is equivalent to an otherwise identical momentarily comoving inertial observer. They found that the consistency can be achieved only in a rather limited neighborhood around the observer with linear dimensions that are negligibly small compared to the characteristic acceleration length of the observer. Furthermore, Nikolić [16] analyzed the rotational motion in special relativity by calculating transformation relations from Nelson’s general transformation relations [2, 3].
@@
In the present work, we investigate the coordinate transformation relations for a non-uniform acceleration such as a frame executing simple harmonic motion (SHM) and subsequently the transformation relations for a frame having uniform circular motion (UCM). However, in special relativity, when the motion of a test particle is subject to a restoring force along the $X$-direction, the mass of the particle can not be considered as constant. But, it changes according to the law of special theory of relativity to which MacColl [17] calculated a general expression of the time period of the relativistic oscillation, which is a function of the total energy of the system. Such relativistic harmonic oscillation can be used to admit a consistent procedure for the quantization of the system [18]. However, the formal investigation on the harmonic oscillator in relativity using the Lagrange formalism [19] and the Hamiltonian formalism [20] indicates that even though the term “harmonic oscillator” has been generally used, but the results obtained during the calculations are no longer harmonic in nature. Instead, it is “anharmonic” in nature. It indicates that if the maximum velocity of the SHM is comparable to the speed of light, then the harmonic oscillation becomes an anharmonic in nature. The nature of the relativistic harmonic oscillation and the associated anharmonicity due to the relativistic effect have been experimentally studied using the ultracold lithium atoms in the third band of an optical lattice and reported by Fujiwara et al. [21]. Here, it has been stated that the measured worldline shapes, relativistic anharmonicity, monopole oscillations, and relativistic dephasing of oscillator ensembles are in good agreement with the theoretical predictions of the relativistic harmonic oscillation. It is to be mentioned that, although the simple harmonic oscillation has been studied extensively in the regime of relativity, but only a few works on relativistic coordinate transformation between an inertial frame and a frame executing SHM have been reported in literature to our knowledge. For example, Perepelkin et al. [22] discussed the transformation relation for a frame having velocity which is a sinusoidal function of time by using generalized Galilean transformation. Therefore, it would be very much important and appropriate, to study the transformation relations for the frame undergoing SHM through a more general procedure. This motivated us to calculate such coordinate transformation relations in the present work.
@@
So in the present context, in order to carry out the calculations to obtain the coordinate transformations between an inertial frame and a frame executing SHM, we have to first derive Rindler like transformation equations for coordinates frames having non-uniform acceleration. Since the circular motion of an object can be described by two SHMs with same amplitudes in perpendicular directions having phase difference of $\pi/2$ radians, so the transformation relations for the frame executing SHM can be used to understand the dynamics of a moving frame in a circular path. As a result, we obtain the coordinate transformation relations for a frame of reference moving in a circular path. The results in the present work will have immense applications in the study of the test particles in the non-uniformly accelerated frames such as frames executing SHM and the frames moving in a circular path.
@@
We organized our paper as: In Section 2 we obtain the transformation relations for a frame of reference executing simple harmonic motion (SHM) along $X$-direction. Additionally, we also calculated the similar transformation relations for the frame of reference executing SHM along the $Y$-direction with initial phase of $\pi/2$ radians. And then in Section 3, as mentioned above since the circular motion can be described by using the two SHMs in perpendicular direction and phase difference of $\pi/2$ radians, towards the end we obtained the transformation relations for a frame that is moving in a circular path. In Section 4 we compared our results with the existing coordinate transformation relations for the frames executing SHM and UCM available in literature. At last in Section 5 we discussed the results obtained in the above calculations and concluded our findings.
@@ 2 Transformation relations for a frame of reference executing SHM
The original Rindler transformation equation is meant for constant acceleration. However, in the present work, we shall derive the transformation relations in which proper acceleration or coordinate acceleration is a function of time. Here, we consider two inertial frames $S(ct,x,y,z)$ and $S^{\prime}(ct^{\prime},x^{\prime},y^{\prime},z^{\prime})$, where the $S^{\prime}$ frame is moving with uniform velocity $v$ with respect to $S$ frame along the $X$-direction. Then the coordinate velocity and acceleration of a particle along the $X$-direction in the $S$ frame can be expressed as $u_{x}=\frac{dx}{dt}$ and $a_{x}=\frac{du_{x}}{dt}$, respectively. Similarly, the coordinate velocity and acceleration of a particle along the $X$-direction in the $S^{\prime}$ frame can be expressed as $u^{\prime}_{x}=\frac{dx^{\prime}}{dt^{\prime}}$ and $a^{\prime}_{x}=\frac{du^{\prime}_{x}}{dt^{\prime}}$, respectively. As available in the literature [23, 24, 25], the proper time is the time measured by a clock tied to the frame $S^{\prime}$ given by $\tau=\int\sqrt{1-v^{2}/c^{2}}dt$. Furthermore, the proper velocity and proper acceleration are defined as $dx/d\tau$ and $d(dx/d\tau)/dt$ respectively. We denote the proper acceleration by $\alpha$ in our work. Physically, proper acceleration is that which is measured by an accelerometer carried along with the accelerated object’s frame. Most specifically for a frame executing simple harmonic motion (SHM) the coordinate acceleration is a sinusoidal function of coordinate time in that frame. Below, we derive the transformation relations for the case when the frame is oscillating along the $X$-direction having displacement proportional to some sinusoidal function of angular frequency $\omega$ as defined in the moving frame $S^{\prime}$.
@@
where $u_{x}=\frac{dx}{dt}$ and $a_{x}=\frac{du_{x}}{dt}$ respectively are the coordinate velocity and the coordinate acceleration in the $S$ frame; $a^{\prime}_{x}=\frac{du^{\prime}_{x}}{dt^{\prime}}=\frac{d^{2}x^{\prime}}{dt^{\prime 2}}$ represents the coordinate acceleration of a particle in the $S^{\prime}$ frame. Therefore, for a particle that has acceleration $a^{\prime}_{x}$ with respect to an inertial frame $S^{\prime}$ which instantaneously accompanies the particle $(u_{x}=v)$, we can write:
@@
In case of Castillo & Sanchez [8] they assumed that the acceleration $a^{\prime}_{x}$ is constant and obtained the Rindler’s transformation equation [26]. However, in our present case we shall consider $a^{\prime}_{x}$ to be a function of $(x^{\prime},t^{\prime})$ and proceed as follows.
@@
The quantity $u_{x}$ in the right hand side is actually representing $v$. So we can identify $\left(1-\frac{u_{x}^{2}}{c^{2}}\right)^{1/2}dt=d\tau$, where $\tau$ is the proper time. Additionally from equation (2) we can also find, the acceleration $a_{x}^{\prime}$ becomes equal to the proper acceleration ($\alpha$), because we have taken $u_{x}=v$. The details are available in [25]. This is because $a_{x}^{\prime}$ is actually equivalent to the differential co-efficient of proper velocity with respect to co-ordinate time. Also, since the proper time $\tau$ is inherently related to the coordinate time $t^{\prime}$ in the $S^{\prime}$ frame, so in the present context we can easily express the acceleration $a^{\prime}_{x}$ as a function of ($\tau,x^{\prime}$). Thus, we can finally write:
@@
Integrating on both sides,
@@
Equations (6) and (7) with appropriate constants of integration will help us to find appropriate co-ordinate transformation equations where the acceleration is not constant, but varying with time $(t^{\prime})$.
So, in our present case, we take $a^{\prime}_{x}$ to be a function of time $(t^{\prime})$ and we shall consider two standard cases: (i) the frame $S^{\prime}$ undergoing Simple Harmonic Motion (SHM) and (ii) the frame $S^{\prime}$ undergoing Uniform Circular Motion (UCM).
@@ 2.1 The frame $S^{\prime}$ undergoing SHM in $X$-direction:
In the first case, we consider that acceleration $a^{\prime}_{x}$ is a function of proper time $\tau$ or $(t^{\prime})$ (we note $\tau$ and $t^{\prime}$ are identical in our present problem) and represents a simple harmonic motion (SHM). So we may take a displacement $x^{\prime}=r_{0}\sin(\omega\tau)$ such that the velocity of the oscillation $u^{\prime}_{x}=\frac{dx^{\prime}}{d\tau}=r_{0}\omega\cos(\omega\tau)$ and the acceleration $a^{\prime}_{x}=\frac{d^{2}x^{\prime}}{d\tau^{2}}=-r_{0}\omega^{2}\sin(\omega\tau)$, where $r_{0}$ is the amplitude of the SHM and $\omega$ is the angular frequency (to note $a^{\prime}_{x}$ is co-ordinate acceleration and $\omega$ is measured in the $S^{\prime}$ frame). Therefore, substituting $a_{x}^{\prime}=-r_{0}\omega^{2}\sin(\omega\tau)$ in equation (4) we obtain:
@@
Now, integrating on both side of the equation (9) we obtain as:
@@
The right hand side (RHS) of the above equation (13) is analytically not integrable. However, as in this equation $r_{0}\omega$ represents the maximum value of the linear velocity of the frame executing simple harmonic motion (SHM), and for all practical purposes we assume this velocity will be much smaller than the speed of light $(c)$. Thus, assuming $\frac{r_{0}\,\omega}{c}<<1$, we can write: $\frac{r_{0}\omega}{c}\cos(\omega\tau)\sim 0$ and hence the right hand side (RHS) of equation (13) can be expanded in Taylor series around zero (which is also called Maclaurin Series) in the power of $r_{0}\omega/c$. Now, expanding the integrand in the RHS of equation (13) and neglecting the terms higher than $(r_{0}\omega/c)^{3}$ in order, we can obtain from the above equation (13) as below:
@@
Again, since as stated above, $r_{0}\omega/c<<1$ for all practical purposes, we have $\frac{r_{0}\omega}{c}\cos(\omega\tau)\sim 0$, and hence, we can take an approximation by neglecting the terms higher than $(r_{0}\omega/c)^{3}$ as before. So as before, expanding the integrand in the RHS of equation (15) in Taylor series and utilizing the above approximation we can obtain from equation (15) as:
@@
Equations (14) and (16) represent the world-line of a particle undergoing time varying acceleration, in terms of its proper time ($\tau$). And using (14) and (16) we can obtain as:
$x^{2}-(ct)^{2}=\mathcal{F}(\tau)$ (17)
where $\mathcal{F}(\tau)=\Big[r_{0}\sin(\omega\tau)+\frac{r_{0}^{3}\omega^{2}}{72c^{2}}\Big\{9\sin(\omega\tau)+\sin(3\omega\tau)\Big\}+K_{2}\Big]^{2}-\Big[c\tau+\frac{r_{0}^{2}\omega}{8c}\Big\{2\omega\tau+\sin(2\omega\tau)\Big\}+cK_{3}\Big]^{2}$. Thus for a given $\tau$, the quantity $a^{\prime}_{x}$ can be treated as constant, so using the expression $a_{x}^{\prime}=-r_{0}\omega^{2}\sin(\omega\tau)$, we can rewrite the function $\mathcal{F}$ as follows:
$\mathcal{F}$ $=\Bigg[-\frac{a^{\prime}_{x}}{\omega^{2}}+\frac{r_{0}^{3}\omega^{2}}{72c^{2}}\left\{-\frac{9a^{\prime}_{x}}{r_{0}\omega^{2}}-4\left(\frac{a^{\prime}_{x}}{r_{0}\omega^{2}}\right)^{3}+3\left(\frac{a^{\prime}_{x}}{r_{0}\omega^{2}}\right)\right\}+K_{2}\Bigg]^{2}$
$-\Bigg[\frac{c}{\omega}\sin^{-1}\left(-\frac{a^{\prime}_{x}}{r_{0}\omega^{2}}\right)\left(1+\frac{r_{0}^{2}\omega^{2}}{4c^{2}}\right)-\frac{r_{0}}{4c}\frac{a^{\prime}_{x}}{\omega}\sqrt{1-\left(\frac{a^{\prime}_{x}}{r_{0}\omega^{2}}\right)^{2}}+cK_{3}\Bigg]^{2}$ (18)
which is a constant for a given $\tau$. So, the world-line of a particle with constant acceleration is a hyperbola [25, 8]. This fact verifies that the calculations carried out in the present work upto this stage, satisfies the limiting conditions of Rindler’s Hyperbola.
Now, as mentioned above, the frame $S^{\prime}$ is an inertial frame which instantaneously accompanies the particle undergoing time varying acceleration and it is evident from the calculations shown in Appendix A that, the frame $S^{\prime}$ is a Fermi-Walker transported frame [24, 1, 14]. So, in this situation, the time ($\tau$) measured by the particle is equal to the coordinate time ($t^{\prime}$) of the $S^{\prime}$ frame.
Therefore, in the present context, as $\tau$ can be the time in the frame executing SHM, so we may replace $\tau$ by $t^{\prime}$, where $t^{\prime}$ represents the time as measured from the accelerated frame (as discussed just before equation (4)). Therefore, substituting $\tau=t^{\prime}$ we can obtain from equations (14) and (16) as follows:
@@
As mentioned above, the frame $S^{\prime}$ is essentially a Fermi-Walker transported frame, and locally coincides with an inertial frame along the observer’s world-line. So, the frame $S^{\prime}$ is locally/ instantaneously inertial and preserve the Minkowski metric. So, the transformation relations (21)–(24) are locally Lorentz-type [25, 24].
@@
Then we integrate both sides of equation (27), with limit 0 to $T$, for the unprimed and 0 to $T^{\prime}$, for the primed frame. i.e.
@@
Comparing (29) and (30), one can say that, we obtained $(1/4)$ as the co-effecient of the second term in RHS, corresponding to $(3/16)$ for the same quantity as obtained by other authors [19, 21, 27].
This fact once again confirms that all our past calculations based on which we finally obtained the time periods (in the above), match with the same set of results obtained by other authors in past.